OpenStudy (anonymous):

i need help with steps understanding how to slove this problem 1,210=1.10x

OpenStudy (saifoo.khan):

x = 1210/1.10

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (anonymous):

1.10x = 1210 x = 1210/1.10 x = 1100

OpenStudy (anonymous):

how did u get that

OpenStudy (saifoo.khan):

Simply by cross multiplying!

OpenStudy (anonymous):

Result = x= 12,1

OpenStudy (anonymous):

10x = 1,210 100x=1210 x= 1210/100 x = 12,1

OpenStudy (anonymous):

1.10x = 1210 1.10x/1.10 = 1210/1.10 x = 1210/1.10 x = 1100